OpenStudy (anonymous):

what is 3s^2+5s+2=0 ?????

OpenStudy (saifoo.khan):

\[s = -1\] \[s = - \frac23\]

hero (hero):

3s^2+5s + 2 = 0 3s^2 + 3s + 2s + 2 = 0 3s(s+1)+2(s+1) = 0 (s+1)(3s+2)=0 s = -1, -2/3

hero (hero):

I hope that helps

OpenStudy (anonymous):

Thanks, does the FOIL method work for that then?

hero (hero):

I'm not a big fan of FOIL, but if you foil (s+1)(3s+2), you should get back to the original trinomial

OpenStudy (anonymous):

thanks a lot

hero (hero):

Looks like Saifoo lost his Mod Status

OpenStudy (saifoo.khan):

i never got a mod status. :p

hero (hero):

You did. You had it, you just never knew it. We were discussing this earlier.